12. The tax multiplier for the below economy would be

domestic output AE, closed economy government
200 230 20
250 270 20
300 310 20
350 350 20
400 390 20
450 430 20
500 470 20

A) -2
B) -3

C) -4
D) -5

13. If the economy is operating at a point above the equilibrium point, then

A) AE is greater than GDP

B) GDP is greater than AE
C) The economy is at a sustainable full employment level
D) Business inventories have been depleted

14. Which of the following is NOT true in the Keynesian aggregate expenditures model?

A) Price level is held constant

B) There is a direct relationship between aggregate expenditures and GDP
C) The wealth effect has no bearing on the aggregate expenditures model

D) All changes in spending are autonomous and do not change the MPC

15. Other things equal, of a change in the tastes of American consumers causes them to purchase more foreign goods at each level of US GDP

A) Unemployment will decrease domestically

B) US GDP will fall
C) Inflation will occur domestically
D) US real GDP will rise
Answer #1

12. Ans: -4

Explanation:

From the table, it is seen that when domestic output increases by 50, the AE increases by 40. So,

MPC = 40 / 50 = 0.8

MPS = 1 - MPC = 1 - 0.8 = 0.2

Tax multiplier = -(MPC / MPS) = -(0.8 / 0.2) = -4

13. Ans: GDP is greater than AE

14. Ans: The wealth effect has no bearing on the aggregate expenditures model

Explanation:

AE depends on disposable income, wealth, expectations, demographics. Thus, The wealth effect has bearing on the aggregate expenditures model.

15. Ans: US GDP will fall

Explanation:

GDP = C + I + G + X - M

When, American consumers purchase more foreign goods , import (M) will increase. If M increases, GDP will fall